Susan is 4 times as old as Paul and 1/2 times as old as Mary . What fractions of Paul's age is Mary ?
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

Susan is 4 times as old as Paul and 1/2 times as old as Mary . What fractions of Paul's age is Mary ?
-------
Equations:
s = 4p
s = (1/2)m
-------
(1/2)m = 4p
----
m = 8p
------
Mary is 8 times as old as Paul
-----
Paul is (1/8)th Mary's age
==================================
Cheers,
Stan H.
==================================

Answer by lwsshak3(11628)

Susan is 4 times as old as Paul and 1/2 times as old as Mary . What fractions of Paul's age is Mary ?
let x=Mary's age
x/2=Susan's age
1/4*x/2=x/8= Paul's age
Mary's age/Paul's age=x/(x/8)=8
Mary is 8 times Paul's age.